Consider the utility function U(x, y) = 3x + y, with MUx = 3 and MUy = 1. a) Is the assumption that more is better satisfied for both goods? b) Does the marginal utility of x diminish, remain constant,or increase as the consumer buys more x? Explain. c) What is MRSx,y ? d) Is MRSx,y diminishing, constant, or increasing as the consumer substitutes x for y along an indifference curve? e) On a graph with x on the horizontal axis and y on the vertical axis,draw a typical indifference curve (it need not be exactly to scale, but it needs to reflect accurately whether there is a diminishing MRSx,y).Also indicate on your graph whether the indifference curve will intersect either or both axes.Label the curve U1. f) On the same graph draw a second indifference curve U2,with U2 > U1.
1
Expert's answer
2020-11-27T10:05:08-0500
"Solution"

a.

"MU_x=\frac{\delta u}{\delta x}=3>0, Hence\ MU_x>0\\
MU_y=\frac{\delta u}{\delta y}=1>0, Hence\ MU_y>0\\"

Since "MU_x" and "MU_y" are both greater than zero, the assumption that more is better satisfied for both goods.


b.

"\frac{\delta MU_x}{\delta x}=0" . Hence marginal utility of "x" remains constant as the consumer buys more "x" .


c.

MRSxy is the amount of "y" consumer is willing to give for marginal unit of "x", so that their utility remains the same.


d.

"MRS_{xy}=-\frac{MU_x}{MU_y}=-\frac{3}{1}=-3"

Hence "\frac{\delta MRS_{xy}}{\delta x}=0 \implies MRS" is constant along an indifference curve.


e.



From the above graph with x on the horizontal axis and y on the vertical axis, we can see that the indifference curve is touching both axis, and as indifference curve is a straight line. Hence MRS is a constant well known as a perfect substitution.


f.



"U_2>U_1"

Both "U_1" and "U_2" are parallel hence they intersect both axes.